### v3

@@ 7.1.1 Improvement on ICL capabilities
Recently, a growing amount of work has focused on enhancing ICL performance under various scenarios. In this section, we trace the development of this field and summarize relevant works.
MIMIC-IT [137] combines in-context learning with instruction tuning by building an instruction dataset formatted with multimodal context. Some other works explore improving few-shot learning performance under specific settings. For example, Link-context learning [138] focuses on the causal relationships between demonstrations and queries, and casts a contrast training scheme by formulating positive and negative image-description pairs. Similarly, Yang et al.  [139, 140] explore different strategies to optimize demonstration configurations (selections or orderings of in-context samples) to achieve better few-shot performance.
@@ 7.1.2 Applications
In terms of applications in multimodality, M-ICL is mainly used in two scenarios: (1) solving various visual reasoning tasks [141] and (2) teaching LLMs to use external tools [142, 143]. The former involves learning from several task-specific examples and generalizing to a new but similar question. In contrast, examples of tool usage are more fine-grained, typically comprising a chain of steps to fulfill the task.
@@ 7.2 Multimodal chain of thought
CoT is “a series of intermediate reasoning steps” [5]. The technique has been proven to be effective in complex reasoning tasks. The main idea is to prompt LLMs to output not only the final answer but also the reasoning process that leads to the answer, resembling the cognitive process of humans.
Inspired by the success in NLP realms, multiple works [144, 145] have proposed to extend the technique to Multimodal CoT (M-CoT). We first introduce different paradigms for acquiring the M-CoT ability. Then, we delineate more specific aspects of M-CoT, including the chain configuration and the pattern.
@@ 7.2.1 Learning paradigms
There are broadly three ways to acquire the M-CoT ability, i.e. through finetuning and training-free few/zero-shot learning.
Intuitively, the finetuning approach often involves curating specific datasets for M-CoT learning. For example, Lu et al.  [100] construct a scientific question-answering dataset ScienceQA with lectures and explanations, which can serve as sources of learning CoT reasoning.
Compared with finetuning, few/zero-shot learning is more computationally efficient. The few-shot learning approach typically requires hand-crafted in-context examples to teach reasoning step by step. In contrast, the zero-shot learning approach directly prompts with designed instructions [144].
@@ 7.2.2 Chain configuration
Structure and length are two critical aspects of the reasoning chains. In terms of structure, current methods can be divided into single-chain [100] and tree-shape methods [146]. Chain length can be categorized into adaptive and pre-defined formations. The former configuration requires LLMs to decide when to halt the reasoning chains [100], while the latter setting stops the chains with a pre-defined length [147].
@@ 7.2.3 Generation patterns
We summarize the relevant works into (1) an infilling-based pattern and (2) a predicting-based pattern. Specifically, the infilling-based pattern demands deducing steps between surrounding context (previous and following steps) to fill the logical gaps [144]. In contrast, the predicting-based pattern requires extending the reasoning chains given conditions such as instructions and previous reasoning history [142].
@@ 7.3.1 Introduction
Inspired by the success of tool-augmented LLMs [148], some researchers have explored the possibilities of invoking external tools or vision foundation models for visual reasoning tasks. Taking LLMs as helpers with different roles, these works build task-specific or general-purpose visual reasoning systems.
Compared with conventional visual reasoning models, these works manifest several good traits: (1) Strong generalization abilities. Equipped with rich open-world knowledge learned from large-scale pretraining, these systems can easily generalize to unseen objects or concepts with remarkable zero/few-shot performance [149]. (2) Emergent abilities. Aided by the strong reasoning abilities of LLMs, these systems can perform complex tasks, e.g. understanding the deep meaning of an image [18]. (3) Better interactivity and control. Traditional models typically allow a limited set of control mechanisms, while LLM-based systems enable finer control in a user-friendly interface (e.g. click and natural language queries) [150].
For this part, we start with introducing different training paradigms employed in the construction of LLM-Aided Visual Reasoning systems. Then, we delve into the primary roles that LLMs play within these systems.
@@ 7.3.2 Training paradigms
Training-free. With abundant prior knowledge stored in pre-trained LLMs, an intuitive and simple way is to freeze pre-trained models and directly prompt LLMs to fulfill various needs. According to the setting, the reasoning systems can be further categorized into few-shot models [142] and zero-shot models [150].
Finetuning. Some works adopt further finetuning to improve the planning abilities with respect to tool usage [90] or to improve localization capabilities [114] of the system. For example, GPT4Tools [90] collects a tool-related instruction dataset to finetune the model.
@@ 7.3.3 Functions
Regarding what roles LLMs exactly play in LLM-Aided Visual Reasoning systems, existing related works are divided into three types:
@@
We delineate how LLMs serve these roles in the following parts.
LLM as a Controller. In this case, LLMs act as a central controller that (1) breaks down a complex task into simpler sub-tasks/steps and (2) assigns these tasks to appropriate tools/modules. Specifically, LLMs are prompted explicitly to output task planning [151] or, more directly, the modules to call [142, 90, 143]. For example, VisProg [143] prompts GPT-3 to output a visual program, where each program line invokes a module to perform a sub-task.
LLM as a Decision Maker. In this case, complex tasks are solved in a multi-round manner, often in an iterative way [152]. Decision-makers often (1) summarize the context to decide whether to finish the task and (2) organize the answer in a user-friendly way.
LLM as a Semantics Refiner. When LLM is used as a Semantics Refiner, researchers mainly utilize its rich linguistic and semantic knowledge. Specifically, LLMs are often instructed to integrate information into fluent natural language sentences [153] or generate texts according to different specific needs [150, 149, 154].
@@ 8 CHALLENGES AND FUTURE DIRECTIONS
  • Current MLLMs are limited in processing multimodal information of long context. This restricts the development of advanced models with more multimodal tokens, e.g. long-video understanding and long documents interleaved with images and text.
  • MLLMs should be upgraded to follow more complicated instructions. For example, a mainstream approach to generating high-quality question-answer pair data is still prompting closed-source GPT-4V because of its advanced instruction-following capabilities, while other models generally fail to achieve such goals.
  • There is still a large space for improvement in techniques like M-ICL and M-CoT. Current research on the two techniques is still rudimentary, and the related capabilities of MLLMs are still weak. Therefore, explorations on the underlying mechanisms and potential improvements are promising.
@@
  • Safety issues. Similar to LLMs, MLLMs can be vulnerable to crafted attacks. In other words, MLLMs can be misled to output biased or undesirable responses. Thus, improving model safety will be an important research topic.
  • Interdisciplinary research. Given the strong generalization capabilities and abundant pre-trained knowledge of MLLMs, a promising research direction could be utilizing MLLMs to boost research fields of natural sciences, e.g. leveraging MLLMs for analysis of medical images or remote sensing images. To achieve this goal, injecting domain-specific multimodal knowledge into MLLMs might be necessary.
@@ 9 CONCLUSION
In this paper, we perform a survey of the existing MLLM literature and offer a broad view of its main directions, including the basic recipe and related extensions. Moreover, we underscore the current research gaps that need to be filled and point out some promising research directions. We hope this survey can offer readers a clear picture of the current progress of MLLM and inspire more relevant works. In light of the fact that the era of MLLM has only just begun, we will keep updating this survey and hope it can inspire more research. An associated GitHub link collecting the latest papers is available here.
FUNDING
This work was supported in part by the National Natural Science Foundation of China (62222213, 62406264, U22B2059, U23A20319, 62072423 and 61727809) and the Young Scientists Fund of the Natural Science Foundation of Sichuan Province (2023NSFSC1402).
AUTHOR CONTRIBUTIONS
C.F. was the project leader. S.Y., C.F. and S.Z. performed the literature review. K.L., X.S., T.X. and E.C. provided insights of the relevant fields. S.Y., C.F. and S.Z. wrote the article. All authors discussed and advised on the manuscript.
Conflict of interest statement. None declared.

### v4

@@ 7.1.1 Improvement on ICL capabilities
Recently, a growing amount of work has focused on enhancing ICL performance under various scenarios. In this section, we trace the development of this field and summarize some relevant works.
MIMIC-IT [172] combines in-context learning with instruction tuning by building an instruction dataset formatted with multimodal context. The model instruction tuned on the introduced dataset shows improved few-shot performance in the caption task. Emu [173] extends the idea of Flamingo [74] by introducing extra modalities in model generation and corresponding training corpus. Aided by the introduced vision decoder, i.e. Stable Diffusion, the model learns from extra vision supervision and supports more flexibility in output format and in-context reasoning. Specifically, apart from answering in pure text, the model can also give responses in the form of images. Sheng et al. [174] adopt a similar idea and try to extend output modalities into both text and image. Instead of adopting a specialized encoder for images, the work adopts a unified quantization scheme with a shared embedding layer.
Some other works explore improving few-shot learning performance under specific settings. Link-context learning [175] focuses on strengthening the causal link between image-label pairs and casts a contrast training scheme by formulating positive and negative image-description pairs. MMICL [176] aims to augment the capabilities in reasoning with multiple related images. To strengthen the link between image and text, the work proposes a context scheme to transform interleaved image-text data into a uniform format. Jeong [177] finds that when inserting a small fraction of incoherent images/text as noise, MLLMs can be misled to give responses inconsistent with the context. Based on the observation, the work accordingly proposes a pre-filtering method to remove irrelevant context and facilitate more coherent responses.
@@ 7.1.2 Applications
In terms of applications in multimodality, M-ICL is mainly used in two scenarios: (1) solving various visual reasoning tasks [178, 74, 22, 179, 180] and (2) teaching LLMs to use external tools [169, 181, 170]. The former usually involves learning from a few task-specific examples and generalizing to a new but similar question. From the information provided in instructions and demonstrations, LLMs get a sense of what the task is doing and what the output template is and finally generate expected answers. In contrast, examples of tool usage are more fine-grained. They typically comprise a chain of steps that could be sequentially executed to fulfill the task. Thus, the second scenario is closely related to CoT (see §7.2).
@@ 7.2 Multimodal Chain of Thought
As the pioneer work [8] points out, CoT is “a series of intermediate reasoning steps”, which has been proven to be effective in complex reasoning tasks [8, 182, 183]. The main idea of CoT is to prompt LLMs to output not only the final answer but also the reasoning process that leads to the answer, resembling the cognitive process of humans.
Inspired by the success in NLP, multiple works [184, 185, 186, 187] have been proposed to extend the unimodal CoT to Multimodal CoT (M-CoT). We first introduce different paradigms for acquiring the M-CoT ability (§7.2.1). Then, we delineate more specific aspects of M-CoT, including the chain configuration (§7.2.2) and the pattern (§7.2.3).
@@ 7.2.1 Learning Paradigms
The learning paradigm is also an aspect worth investigating. There are broadly three ways to acquire the M-CoT ability, i.e. through finetuning and training-free few/zero-shot learning. The sample size requirement for the three ways is in descending order.
Intuitively, the finetuning approach often involves curating specific datasets for M-CoT learning. For example, Lu et al. [116] construct a scientific question-answering dataset ScienceQA with lectures and explanations, which can serve as sources of learning CoT reasoning, and finetune the model on this proposed dataset. Multimodal-CoT [185] also uses the ScienceQA benchmark but generates the output in a two-step fashion, i.e. the rationale (chain of reasoning steps) and the final answer based on the rationale. CoT-PT [187] learns an implicit chain of reasoning through a combination of prompt tuning and step-specific visual bias.
Compared with finetuning, few/zero-shot learning is more computationally efficient. The main difference between them is that the few-shot learning typically requires hand-crafting some in-context examples so that the model can learn to reason step by step more easily. In contrast, the zero-shot learning does not require any specific example for CoT learning. In this case, models learn to use the embedded knowledge and the reasoning ability without explicit guidance by prompting designed instructions like “Let’s think frame by frame” or “What happened between these two keyframes” [184, 186]. Similarly, some works [22, 188] prompt models with descriptions of the task and tool usage to decompose complex tasks into sub-tasks.
@@ 7.2.2 Chain Configuration
Structure and length are two critical aspects of the reasoning chains. In terms of structure, current methods can be divided into single-chain and tree-shape methods. Reasoning with a single chain is a paradigm widely used in various methods [116, 185]. Specifically, the step-by-step reasoning process forms a single question-rationale-answer chain. Recently, some methods have explored using a more complicated scheme, i.e. tree-shape chain, for reasoning. Specifically, DDCoT [189] breaks down a question into multiple sub-questions, each of which is solved by LLM itself or visual experts to generate rationales. Then the LLM aggregates and reasons with the rationales to form the final answer. With respect for chain length, it can be categorized into adaptive and pre-defined formations. The former configuration requires LLMs to decide on their own when to halt the reasoning chains [169, 116, 22, 185, 188, 170], while the latter setting stops the chains with a pre-defined length [187, 79, 186, 184].
@@ 7.2.3 Generation Patterns
How the chain is constructed is a question worth studying. We summarize the current works into (1) an infilling-based pattern and (2) a predicting-based pattern. Specifically, the infilling-based pattern demands deducing steps between surrounding context (previous and following steps) to fill the logical gaps [186, 184]. In contrast, the predicting-based pattern requires extending the reasoning chains given conditions such as instructions and previous reasoning history [169, 116, 22, 185, 188, 170]. The two types of patterns share a requirement that the generated steps should be consistent and correct.
@@ 7.3.1 Introduction
Inspired by the success of tool-augmented LLMs [190, 191, 192, 193], some researches have explored the possibilities of invoking external tools [107, 170, 22, 169] or vision foundation models [22, 79, 80, 194, 188, 195, 196] for visual reasoning tasks. Taking LLMs as helpers with different roles, these works build task-specific [197, 198, 79] or general-purpose [181, 22, 188, 169, 170] visual reasoning systems.
Compared with conventional visual reasoning models [199, 200, 201], these works manifest several good traits: (1) Strong generalization abilities. Equipped with rich open-world knowledge learned from large-scale pretraining, these systems can easily generalize to unseen objects or concepts with remarkable zero/few-shot performance [169, 170, 195, 197, 198, 202]. (2) Emergent abilities. Aided by strong reasoning abilities of LLMs, these systems can perform complex tasks. For example, given an image, MM-REACT [22] can interpret the meaning beneath the surface, such as explaining why a meme is funny. (3) Better interactivity and control. Traditional models typically allow a limited set of control mechanisms and often entail expensive curated datasets [203, 204]. In contrast, LLM-based systems have the ability to make fine control in a user-friendly interface (e.g. click and natural language queries) [79].
For this part, we start with introducing different training paradigms employed in the construction of LLM-Aided Visual Reasoning systems (§7.3.2). Then, we delve into the primary roles that LLMs play within these systems (§7.3.3).
@@ 7.3.2 Training Paradigms
Training-free. With abundant prior knowledge stored in pre-trained LLMs, an intuitive and simple way is to freeze pre-trained models and directly prompt LLMs to fulfill various needs. According to the setting, the reasoning systems can be further categorized into few-shot models [169, 181, 22, 170] and zero-shot models [197, 79]. The few-shot models entail a few hand-crafted in-context samples (see §7.1) to guide LLMs to generate a program or a sequence of execution steps. These programs or execution steps serve as instructions for corresponding foundation models or external tools/modules. The zero-shot models take a step further by directly utilizing LLMs’ linguistics/semantics knowledge or reasoning abilities. For example, PointCLIP V2 [197] prompts GPT-3 to generate descriptions with 3D-related semantics for better alignment with corresponding images. In CAT [79], LLMs are instructed to refine the captions according to user queries.
Finetuning. Some works adopt further finetuning to improve the planning abilities with respect to tool usage [107] or to improve localization capabilities [142, 205] of the system. For example, GPT4Tools [107] introduces the instruction-tuning approach (see §3.2). Accordingly, a new tool-related instruction dataset is collected and used to finetune the model.
@@ 7.3.3 Functions
In order to further inspect what roles LLMs exactly play in LLM-Aided Visual Reasoning systems, existing related works are divided into three types:
@@
The first two roles are related to CoT (see §7.2). It is frequently used because complex tasks need to be broken down into intermediate simpler steps. When LLMs act as controllers, the systems often finish the task in a single round, while multi-round is more common in the case of the decision maker. We delineate how LLMs serve these roles in the following parts.
LLM as a Controller. In this case, LLMs act as a central controller that (1) breaks down a complex task into simpler sub-tasks/steps and (2) assigns these tasks to appropriate tools/modules. The first step is often finished by leveraging the CoT ability of LLMs. Specifically, LLMs are prompted explicitly to output task planning [181] or, more directly, the modules to call [169, 107, 170]. For example, VisProg [170] prompts GPT-3 to output a visual program, where each program line invokes a module to perform a sub-task. In addition, LLMs are required to output argument names for the module input. To handle these complex requirements, some hand-crafted in-context examples are used as references [169, 181, 170]. This is closely related to the optimization of reasoning chains (see §7.2), or more specifically, the least-to-most prompting [206] technique. In this way, complex problems are broken down into sub-problems that are solved sequentially.
LLM as a Decision Maker. In this case, complex tasks are solved in a multi-round manner, often in an iterative way [195]. Decision-makers often fulfill the following responsibilities: (1) Summarize the current context and the history information, and decide if the information available at the current step is sufficient to answer the question or complete the task; (2) Organize and summarize the answer to present it in a user-friendly way.
LLM as a Semantics Refiner. When LLM is used as a Semantics Refiner, researchers mainly utilize its rich linguistics and semantics knowledge. Specifically, LLMs are often instructed to integrate information into consistent and fluent natural language sentences [202] or generate texts according to different specific needs [79, 197, 198].
@@ 8 Challenges and Future Directions
  • Current MLLMs are limited in processing multimodal information of long context. This restricts the development of advanced models with more multimodal tokens, e.g. long-video understanding, and long documents interleaved with images and text.
  • MLLMs should be upgraded to follow more complicated instructions. For example, a mainstream approach to generating high-quality question-answer pair data is still prompting closed-source GPT-4V because of its advanced instruction-following capabilities, while other models generally fail to achieve.
  • There is still a large space for improvement in techniques like M-ICL and M-CoT. Current research on the two techniques is still rudimentary, and the related capabilities of MLLMs are weak. Thus, explorations of the underlying mechanisms and potential improvement are promising.
@@
  • Safety issues. Similar to LLMs, MLLMs can be vulnerable to crafted attacks [177, 207, 208]. In other words, MLLMs can be misled to output biased or undesirable responses. Thus, improving model safety will be an important topic.
@@ 9 Conclusion
In this paper, we perform a survey of the existing MLLM literature and offer a broad view of its main directions, including the basic recipe and related extensions. Moreover, we underscore the current research gaps that need to be filled and point out some promising research directions. We hope this survey can offer readers a clear picture of the current progress of MLLM and inspire more works.